# A Mathematical Theory for Learning Semantic Languages by Abstract Learners

### v2

@@ III-B Poisson Learner
In this section, we extend the concept of the 1-skill learner to the Poisson learner, where multiple novel skills may be learned from a text but only with a certain probability.
@@
Analogous to the argument for the 1-skill learner, the testing error for a Poisson learner is characterized in (4) with $p$ given in (7).
Figure 4: The plot illustrates the computed values of $p$ as a function of $R$, which ranges from 0.1 to 5 with a step size of 0.1, for each $c$ value (1, 2, 3, 4, 5). The computation is based on the recursive equation $p^{(i+1)}=1-P_{\rm suc}(ce^{-cR(1-p^{(i)})})$ with the initial condition $p^{(0)}=1$, where $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$.
In Figure 4, we illustrate the effect of the parameters $R$ and $c$ on the values of $p$ for the 2-skill learner. We compute $p$ by using the recursive equation $p^{(i+1)}=1-P_{\rm suc}(ce^{-cR(1-{p^{(i)}})})$ with the initial condition $p^{(0)}=1$, where $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$. The number of iterations is set at 10,000. A percolation threshold, indicated by a sudden change in the values of $p$, exists for $c=4$ (respectively, 5) when $R$ is approximately 0.5 (resp. 0.6). The percolation thresholds for the 2-skill learner are significantly lower than those for the 1-skill learner, demonstrating that the 2-skill learner is a much faster learner than the 1-skill learner.
@@ IV Learning the association of skills
When the training is complete, a learner can not only learn a fraction of skills but also learn the associations between these learned skills. We say two learned skills are associated if they appear in the same text. As a learned skill can determine whether it is in a text, we can generate a skill association graph for learned skills by adding an edge between two learned skill nodes if they appear in the same text. For instance, the two learned skills, $a$ and $b$, shown in Figure 2, are associated as they appear in the same text. After training, we can generate a skill association graph with nodes $a$ and $b$, and an edge connecting these two nodes.
Knowing the structure of the skill association graph is crucial, as it can be utilized for inference purposes, such as predicting the next skill given a set of skills. This is analogous to giving an LLM a prompt in the form of a text with a set of skills and asking the LLM to predict the next skill and generate a text based on the predicted skill. An essential question arises: Are most of the learned skills interconnected in a way that facilitates their use for prediction? This question guides us to perform a giant component analysis of the skill association graph. Note that a giant component in a random graph is a connected subgraph whose size is proportional to the size of the graph. It is well-known [9] that there exists at most one giant component in a random graph.
Our approach to this problem is based on the site (or node) percolation analysis as described in [9]. As in Section III-B, we consider the skill-text bipartite graph with $|S|$ skill nodes and $D$ text nodes. It is important to note that not every skill is learned upon the completion of training. As previously shown, the probability $p_{s}$ that a randomly selected skill node is learned is equal to $1-e^{-cR(1-p)}$, i.e.,
$p_{s}=1-e^{-cR(1-p)},$ (8)
@@
To apply the site percolation analysis for the skill-text bipartite graph, we assume that skill nodes are learned independently with probability $p_{s}$. Let $\mu_{s}$ be the probability that a skill node is not connected to the giant component via one of its edges. Also let $\mu_{t}$ be the probability that a text node is not connected to the giant component via one of its edges. Recall from Proposition 3 that the degree distribution of the text nodes (respectively, skill nodes) is Poisson with mean $c$ (respectively, $cR$). Note that a skill node cannot reach the giant component via one of its edges, called the tagged edge, if all the excess edge of the text node connected to the tagged edge do not lead to the giant component. Since the excess degree distribution of a Poisson degree distribution is also Poisson with the same mean, we then have
$\mu_{s}=\sum_{k=0}^{\infty}e^{-c}\frac{c^{k}}{k!}(\mu_{t})^{k}=e^{-c(1-\mu_{t})}.$ (9)
Now we derive the governing equation for $\mu_{t}$. If the skill end of a randomly selected edge is not learned, one cannot reach the giant component via this edge, which occurs with probability $1-p_{s}$. Conversely, if the skill end of a randomly selected edge (the tagged edge) is learned, then one cannot reach the giant component if all the excess edges of the skill node connected to the tagged edge do not lead to the giant component. Thus, we have
$\mu_{t}$ $=$ $(1-p_{s})+p_{s}\sum_{k=0}^{\infty}e^{-cR}\frac{{(cR)}^{k}}{k!}(\mu_{s})^{k}$ (10)
$=$ $(1-p_{s})+p_{s}e^{-Rc(1-\mu_{s})}.$
@@
$\mu_{t}^{(i+1)}=(1-p_{s})+p_{s}e^{-Rc(1-\mu_{s}^{(i)})},$ (11)
with the initial condition $\mu_{s}^{(0)}=\mu_{t}^{(0)}=0$. Note that a randomly selected skill node is in the giant component if the skill node is learned and at least one of its edges can reach the giant component. Thus, the probability that a randomly selected skill node is in the giant component, denote by $p_{G}$, is
$p_{s}(1-\sum_{k=0}^{\infty}e^{-cR}\frac{{(cR)}^{k}}{k!}(\mu_{s})^{k})=p_{s}(1-e^{-Rc(1-\mu_{s})}).$ (12)
Figure 5: The plot of the giant component size $p_{G}$ for the 1-skill learner as a function of $R$ for various values of $c$.
In Figure 5, we present a plot of the giant component size $p_{G}$ as a function of $R$ for the 1-skill learner with values of $c=1,2,3,4,5$. The solid line represents the theoretical value, and the dashed line represents the simulated value obtained by averaging over 100 random bipartite graphs with a total of 20,000 skill and text nodes. This figure offers an intriguing comparison to the testing errors shown in Figure 3. For $c=1$, as indicated in Figure 3, the testing errors decrease with an increase in $R$. However, the learned skills do not form a giant component until $R$ surpasses a threshold, approximately at 1.5, as depicted in Figure 5. While it is relatively easy for the 1-skill learner to acquire skills when the average number of skills in a sampled text is small, it is also challenging to learn the association between two learned skills. Conversely, for $c=5$, learning individual skills is difficult. Yet, when $R$ exceeds a threshold, roughly at 2.5, we observe a significant reduction in testing errors in Figure 3. Additionally, surpassing this same threshold in Figure 5 correlates with the emergence of a giant component with the size near the size of the skill nodes. This suggests that the acquisition of skills and their associations occur simultaneously once this threshold is exceeded. It is noteworthy that discrepancies exist between the theoretical and simulation results, as seen in Figure 5 for $c=1$ and $2$. These discrepancies arise because the percolation analysis assumes skill nodes are learned independently–a premise that does not hold when $c$ is small.
@@
$g_{1}(\mu_{s})=e^{-cp_{s}(1-e^{-Rc(1-\mu_{s})})}.$ (14)
Note that $g_{1}(\mu_{s})$ is a convex function of $\mu_{s}$ when $0\leq\mu_{s}\leq 1$. This can be easily verified by showing the second derivative of $g_{1}(\mu_{s})$ is nonnegative for $0\leq\mu_{s}\leq 1$. Clearly, $\mu_{s}=1$ is the trivial solution of (13).
@@
By the convexity and the fact that $g_{1}(0)>0$, we have a non-trivial solution for some $\mu_{s}<1$ if the slope of $g_{1}(\mu_{s})$ at $\mu_{s}=1$ is greater than the slope of the line $y=\mu_{s}$, i.e.,
${\frac{d}{d\mu_{s}}g_{1}(\mu_{s})}\bigg{|}_{\mu_{s}=1}>1.$ (15)
See an illustration of (15) in Figure 6, where a non-trivial solution is found at the point where the curve $y=g_{1}(\mu_{s})$ intercepts the line $y=\mu_{s}$, given that $c=3$. A straightforward algebra show that the condition in (15) is equivalent to
$c^{2}Rp_{s}>1.$ (16)
Using (8) yields
$c^{2}R(1-e^{-cR(1-p)})>1$ (17)
where $p$ is the probability that the text end of a randomly selected edge is not learned and it is the solution of (7).
For each pair of $c$ and $R$, we can check whether the condition in (17) is satisfied. If it is satisfied, then there exists a giant component. In Figure 7, we show the region (marked in red) where a giant component emerges. On the other hand, the region (marked in blue) does not exist a giant component. As shown in Figure 7, it is observed that $R$ must be sufficiently large for both scenarios where $c$ is large and $c$ is small. Specifically, when $c$ is large (e.g., $c>2$), a large $R$ is necessary to surpass the percolation threshold for $p$ in (7). Surpassing this threshold leads to the learning of a large number of skills, thereby facilitating the automatic emergence of a giant component within the skill association graph due to a large number of skills in a text. Conversely, when $c$ is small (e.g., $c<1$), nearly every skill in a text is learned by the 1-skill learner (since the average number of skills per text is less than 1). This results in a very small $p$. Nonetheless, even in this scenario, a large $R$ is required to learn a sufficient number of skills to meet the threshold for a giant component’s existence as stipulated in (17). We also note from Figure 7 that the minimum $R$ to have a giant component is 1.1 and this is achieved when $c=1.29$.
Figure 7: The region of the slope of the function $g_{1}(\mu_{s})$ at $\mu_{s}=1$ is categorized by color: the red region, where the slope is greater than 1, indicates the existence of a giant component; conversely, the blue region, where the slope is less than 1, signifies the absence of a giant component.
@@ V Hierarchy of skills
One common approach to train a domain-specific LLM is to adopt a pre-trained model, commonly referred to as a foundation model or a basic model, and fine tune it with additional domain-specific texts. Motivated by this, we consider the setting where there are two classes of skills: basic skills and domain-specific skills (see Figure 8 for an illustration). As in Section III-A, we first train a basic model (foundation model) by using a 1-skill learner. As described in Section III-A, we let $p_{s}$ be the probability that a random selected basic skill is learned. Recall that $p_{s}=1-e^{-cR(1-p)}$, where $p$ is the solution of (3).
@@
Now we use the basic model to fine tune a domain-specific model. We assume that there are $|S_{f}|$ domain-specific skills, and a prerequisite of learning a domain-specific skill requires learning a random number of basic skills. Thus, there is a hierarchy of skills. Suppose that the basic skills required by a domain-specific skill are uniformly selected from the set of basic skills $S$. Let $\delta_{\ell}$ be the probability of the number of basic skills required by a domain-specific skill is $\ell$ and $\Lambda_{f}(x)=\sum_{\ell=0}^{\infty}\delta_{\ell}x^{\ell}$ be the corresponding generating function. Then we say a domain-specific skill is learnable if all its basic skills are learned in the basic model, and this occurs with probability
$\sum_{\ell=0}^{\infty}\delta_{\ell}{p_{s}}^{\ell}=\Lambda_{f}(p_{s}).$ (18)
As in the basic model, we collect a subset of $D_{f}$ domain-specific texts and present these $D_{f}$ texts to learn domain-specific skills (as a way of fine tuning the basic model). Let $R_{f}=D_{f}/|S_{f}|$ be the ratio of the number of training domain-specific texts to the number of domain-specific skills. Similarly to the basic model, we make the following two assumptions about a randomly selected domain-specific text:
(Poisson distribution) For a randomly selected domain-specific text $t\in D_{f}$, the number of domain-specific skills in $t$ follows a Poisson random variable with mean $c_{f}$.
(Uniform distribution) The probability that an edge of the (randomly selected) domain-specific text node $t$ is connected to a domain-specific skill node $s$ is $1/|S_{f}|$.
A domain-specific text node is said to be learnable if all the domain-specific skill nodes connected to it are learnable. Otherwise, it is called a non-learnable domain-specific text node. We assume that nothing can be learned from a non-learnable domain-specific text when it is presented to a learner. From (A1d)-(A2d) and (18), we deduce that the probability a randomly selected domain-specific text is learnable is given by
$\sum_{\ell=0}^{\infty}e^{-c_{f}}\frac{c_{f}^{\ell}}{\ell!}(\Lambda_{f}(p_{s}))^{\ell}=e^{-c_{f}(1-\Lambda_{f}(p_{s}))}.$ (19)
Define the bipartite graph formed by the learnable domain-specific skill nodes and the learnable domain-specific text nodes as the learnable domain-specific bipartite graph. In such a bipartite graph, there are approximately $\Lambda_{f}(p_{s})|S_{f}|$ learnable domain-specific skill nodes and $e^{-c_{f}(1-\Lambda_{f}(p_{s}))}D_{f}$ learnable domain-specific text nodes. Furthermore, from Bayes’ formula, we deduce that the degree of a randomly selected learnable domain-specific text node $t$ in the learnable domain-specific bipartite graph follows a Poisson random variable with mean $c_{f}\Lambda_{f}(p_{s})$ as
@@
$=\frac{e^{-c_{f}}\frac{c_{f}^{\ell}}{\ell!}(\Lambda_{f}(p_{s}))^{\ell}}{e^{-c_{f}(1-\Lambda_{f}(p_{s}))}}=e^{-c_{f}\Lambda_{f}(p_{s})}\frac{(c_{f}\Lambda_{f}(p_{s}))^{\ell}}{\ell!}$ (20)
This leads to the following proportion:
@@
Consider the learnable domain-specific bipartite graph.
(Poisson distribution) For a randomly selected learnable domain-specific text $t$, the number of learnable domain-specific skills in $t$ follows a Poisson random variable with mean $c_{f}\Lambda_{f}(p_{s})$.
(Uniform distribution) The probability that an edge of the (randomly selected) learnable domain-specific text node $t$ is connected to a learnable domain-specific skill node $s$ is uniform.
In light of Proposition 6, we can apply the same density evolution analysis used in Section III to analyze the learnable domain-specific bipartite graph. Suppose that we still use the 1-skill learner for fine tuning. Let $p_{f}$ is the probability that the text end of a randomly selected edge in the learnable domain-specific bipartite graph is not learned after fine tuning. Replacing $c$ with $c_{f}\Lambda_{f}(p_{s})$ and $R$ with $R_{f}e^{-c_{f}(1-\Lambda_{f}(p_{s}))}/\Lambda_{f}(p_{s})$ in (3) yields
$p_{f}=1-\exp\Big{(}{-c_{f}\Lambda_{f}(p_{s})e^{-(1-p_{f})c_{f}R_{f}e^{-c_{f}(1-\Lambda_{f}(p_{s}))}}}\Big{)}.$ (21)
Now we derive the probability of testing errors of a randomly selected text from the domain-specific texts. Denote such a probability as $p_{t}$. The test errors consists of two cases: (i) a randomly selected text from the domain-specific texts is not learnable, and (ii) a randomly selected text from the domain-specific texts is learnable but one of the domain-specific skills in that text is not learned. As shown in (19), the probability that a randomly selected text from the domain-specific texts is not learnable is $1-e^{-c_{f}(1-\Lambda_{f}(p_{s}))}$. Since the degree distribution of a domain-specific text in the learnable domain-specific bipartite graph is Poisson with mean $c_{f}\Lambda_{f}(p_{s})$ in (V), the probability that one of the domain-specific skills in a learnable domain-specific text is not learned is
$1-\sum^{\infty}_{k=0}e^{-c_{f}\Lambda_{f}(p_{s})}\frac{(c_{f}\Lambda_{f}(p_{s}))^{k}}{k!}(1-p_{f})^{k}$ (22)
$=$ $1-e^{-c_{f}\Lambda_{f}(p_{s})p_{f}}.$
@@
$(1-e^{-c_{f}\Lambda_{f}(p_{s})p_{f}})e^{-c_{f}(1-\Lambda_{f}(p_{s}))}.$ (23)
Combining Case (i) in (22) and Case (ii) in (23) leads to
$p_{t}$ $=$ $1-e^{-c_{f}(1-\Lambda_{f}(p_{s}))}$ (24)
$+$ $(1-e^{-c_{f}\Lambda_{f}(p_{s})p_{f}})e^{-c_{f}(1-\Lambda_{f}(p_{s}))}$
$=$ $1-e^{-c_{f}(1-\Lambda_{f}(p_{s})(1-p_{f}))}.$
@@
Figure 9: The contour plot of the probability of the test errors $p_{t}$ for the 1-skill learner as a function of $R$ and $R_{f}$ with values $c=3$ and $c_{f}=3$.

### v3

@@ III-B Poisson Learner
In this section, we extend the concept of the 1-skill learner to the Poisson learner, where multiple novel skills may be learned from a text with a certain probability.
@@
Then, analogous to the argument for the 1-skill learner, the testing error for a Poisson learner is characterized in (5) with $p$ given in (8).
Figure 6: The plot illustrates the probability of the testing error $\epsilon$ as a function of $R$ for the 2-skill learner.
In Figure 6, we illustrate the effect of the parameters $R$ and $c$ on the probability of the testing error $\epsilon$ for the 2-skill learner. We first compute $p$ by using the recursive equation $p^{(i+1)}=1-P_{\rm suc}(ce^{-cR(1-{p^{(i)}})})$ with the initial condition $p^{(0)}=1$ and $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$. The number of iterations is set at 10,000. A percolation threshold, indicated by a sudden change in the values of $\epsilon$, exists for $c=4$ (respectively, 5) when $R$ is approximately 0.5 (resp. 0.6). The percolation thresholds for the 2-skill learner are significantly lower than those for the 1-skill learner, demonstrating that the 2-skill learner is a much faster learner than the 1-skill learner.
@@ IV Learning the association of skills
When the training is complete, a learner can learn not only a fraction of skills but also the associations between these learned skills. We say two learned skills are associated if they appear in the same text. Since a learned skill can be identified whenever it is present in a text, we can generate a skill association graph for learned skills by adding an edge between two learned skill nodes if they appear in the same text. For instance, the two learned skills, $a$ and $b$, shown in Figure 2, are associated as they appear in the same text. After training, we can generate a skill association graph with nodes $a$ and $b$, and an edge connecting these two nodes.
Knowing the structure of the skill association graph is crucial, as it can be utilized for inference purposes, such as predicting the next skill given a set of skills. This is analogous to giving an LLM a prompt in the form of a text with a set of skills and asking the LLM to predict the next skill and generate a text based on the predicted skill. An essential question arises: Are most of the learned skills interconnected in a way that facilitates their use for prediction? This question guides us to perform a giant component analysis of the skill association graph. Note that a giant component in a random graph is a connected subgraph whose size is proportional to the size of the graph. It is well-known [9] that there exists at most one giant component in a random graph. The rest of the components are called small components. One important property of a random graph is that small components are trees with high probability.
Our approach to this problem is based on the site (or node) percolation analysis as described in [9]. As in Section III-B, we consider the skill-text bipartite graph with the set of skill nodes ${\cal S}$ and the set of text nodes ${\cal D}$. It is important to note that not every skill is learned upon the completion of training. In fact, as shown in (4), the probability that a randomly selected skill node is learned is given by
$\zeta=1-e^{-cR(1-p)},$ (9)
@@
To apply the site percolation analysis for the skill-text bipartite graph, we assume that skill nodes are learned independently with probability $\zeta$. Let $\mu_{s}$ be the probability that a skill node is connected to a small component via one of its edges. Also, let $\mu_{t}$ be the probability that a text node is connected to a small component via one of its edges. Recall from Proposition 3 that the degree distribution of the text nodes (respectively, skill nodes) is Poisson with mean $c$ (respectively, $cR$). Note that a skill node is only connected to a small component via one of its edges (called the tagged edge) if all the excess edges of the text end of the tagged edge are only connected to small components. Since the excess degree distribution of a Poisson degree distribution is also Poisson with the same mean, we then have (from the tree property for small components) that
$\mu_{s}=\sum_{k=0}^{\infty}e^{-c}\frac{c^{k}}{k!}\mu_{t}^{k}=e^{-c(1-\mu_{t})}.$ (10)
Now we derive the governing equation for $\mu_{t}$. If the skill end of an edge (called the tagged edge) is not learned, then there are no further connected skill nodes via this edge, which occurs with probability $1-\zeta$. Conversely, if the skill end of the tagged edge is learned, then it is only connected to a small component if all the excess edges of the skill end of the tagged edge are only connected to small components. Thus, we have (from the tree property for small components) that
$\mu_{t}$ $=$ $(1-\zeta)+\zeta\sum_{k=0}^{\infty}e^{-cR}\frac{{(cR)}^{k}}{k!}\mu_{s}^{k}$ (11)
$=$ $(1-\zeta)+\zeta e^{-Rc(1-\mu_{s})}.$
@@
$\mu_{t}^{(i+1)}=(1-\zeta)+\zeta e^{-Rc(1-\mu_{s}^{(i)})},$ (12)
with the initial condition $\mu_{s}^{(0)}=\mu_{t}^{(0)}=0$. By induction, it is easy to see from (IV) that $0\leq\mu_{s}^{(i)}<\mu_{s}^{(i+1)}\leq 1$ and $0\leq\mu_{t}^{(i)}<\mu_{t}^{(i+1)}\leq 1$ for all $i$. Thus, the recursion in (IV) converges to a solution in $[0,1]\times[0,1]$ as $i\to\infty$.
Note that a randomly selected skill node is in the giant component if the skill node is learned and at least one of its edges is not connected to a small component. Thus, the probability that a randomly selected skill node is in the giant component, denote by $p_{G}$, is
$p_{G}={\zeta}\Big{(}1-\sum_{k=0}^{\infty}e^{-cR}\frac{{(cR)}^{k}}{k!}\mu_{s}^{k}\Big{)}=\zeta(1-e^{-Rc(1-\mu_{s})}).$ (13)
Figure 7: The plot of the probability $p_{G}$ that a randomly selected skill is in the giant component for the 1-skill learner as a function of $R$ for various values of $c$.
In Figure 7, we present a plot of the probability $p_{G}$ that a randomly selected skill is in the giant component as a function of $R$ for the 1-skill learner with values of $c=1,2,3,4,5$. The solid line represents the theoretical value, and the dashed line represents the simulated value obtained by averaging over 100 random bipartite graphs with a total of 20,000 skill and text nodes. This figure offers an intriguing comparison to the testing errors shown in Figure 5. For $c=1$, as indicated in Figure 5, the testing errors decrease with an increase in $R$. However, the learned skills do not form a giant component until $R$ surpasses a threshold, approximately at 1.5, as depicted in Figure 7. While it is relatively easy for the 1-skill learner to acquire skills when the average number of skills in a sampled text is small, it is also challenging to learn the association between two learned skills. Conversely, for $c=5$, learning individual skills is difficult. Yet, when $R$ exceeds a threshold, roughly at 2.5, we observe a significant reduction in testing errors in Figure 5. Additionally, surpassing this same threshold in Figure 7 correlates with the emergence of a giant component with the size near the size of the skill nodes. This suggests that the acquisition of skills and their associations occur simultaneously once this threshold is exceeded. It is noteworthy that discrepancies exist between the theoretical and simulation results, as seen in Figure 7 for $c=1$ and $2$. These discrepancies arise because the percolation analysis assumes skill nodes are learned independently–a premise that does not hold when $c$ is small.
@@
$g_{1}(\mu_{s})=e^{-c\zeta(1-e^{-Rc(1-\mu_{s})})}.$ (15)
Note that $g_{1}(\mu_{s})$ is a convex function of $\mu_{s}$ when $0\leq\mu_{s}\leq 1$. This can be verified by showing that the second derivative of $g_{1}(\mu_{s})$ is nonnegative for $0\leq\mu_{s}\leq 1$. Clearly, $\mu_{s}=1$ is the trivial solution of (14). By the convexity and the fact that $g_{1}(0)>0$, we have a non-trivial solution for some $\mu_{s}<1$ if the slope of $g_{1}(\mu_{s})$ at $\mu_{s}=1$ is greater than the slope of the line $y=\mu_{s}$, i.e.,
$\left.{\frac{d}{d\mu_{s}}g_{1}(\mu_{s})}\right|_{\mu_{s}=1}>1.$ (16)
This condition is equivalent to
$c^{2}R\zeta>1$ (17)
or, by the representation of $\zeta$ in (9),
$c^{2}R(1-e^{-cR(1-p)})>1.$ (18)
Here, $p$ is the probability that the text end of a randomly selected edge is not learned and is given by the solutions of (3) (respectively, (8)) for a 1-skill learner (respectively, Poisson learner). In Figure 8, we provide an illustration of the condition in (16) for $c=3$ and $R=0.5,1,1.5$ by considering a 1-skill learner. We can see that a non-trivial solution exists at the intersection between the curve $y=g_{1}(\mu_{s})$ and the line $y=\mu_{s}$ for all these cases.
@@
Figure 9: Regions based on the slope of the function $g_{1}(\mu_{s})$ at $\mu_{s}=1$. The red region corresponds to $g_{1}(\mu_{s})>1$, indicating the existence of a giant component, and the blue region corresponds to $g_{1}(\mu_{s})<1$, signifying the absence of a giant component.
For each pair of $c$ and $R$, we can check whether the condition in (18) is satisfied. If it is satisfied, then there exists a non-trivial solution for $\mu_{s}$, implying the existence of a giant component. In Figure 9, we show the region (marked in red) where a giant component emerges. On the other hand, the region (marked in blue) does not exist a giant component. As shown in Figure 9, it is observed that $R$ must be sufficiently large for both scenarios where $c$ is large and $c$ is small. Specifically, when $c$ is large (e.g., $c>2$), a large $R$ is necessary to surpass the percolation threshold for $p$ in (3). Surpassing this threshold leads to the learning of a large number of skills, thereby facilitating the automatic emergence of a giant component within the skill association graph due to a large number of skills in a text. Conversely, when $c$ is small (e.g., $c<1$), nearly every skill in a text is learned by the 1-skill learner (since the average number of skills per text is less than 1). This results in a very small $p$. Nonetheless, even in this scenario, a large $R$ is required to learn a sufficient number of skills to meet the threshold for a giant component’s existence as stipulated in (18). We also note from Figure 9 that the minimum $R$ to have a giant component is 1.1 and this is achieved when $c=1.29$.
@@ V Hierarchy of skills
One common approach to train a domain-specific LLM is to adopt a pre-trained model, commonly referred to as a foundation model or a basic model, and fine tune it with additional domain-specific texts. Motivated by this, we consider the setting where there are two classes of skills: the class of basic skills ${\cal S}$ and the class of domain-specific skills ${\cal S}_{f}$ (see Figure 10 for an illustration). The original set of texts ${\cal T}$ is treated as a basic set of texts, which requires only basic skills to understand. In addition, we also consider an additional set of domain-specific texts ${\cal T}_{f}$, where each requires learning a certain number of domain-specific skills to understand. A prerequisite of learning a domain-specific skill requires learning a random number of basic skills first. Thus, there is a hierarchy of skills.
@@
In particular, following the procedure in Section III-A, we first train a basic model (foundation model) for a 1-skill learner using a set of basic texts ${\cal D}$ sampled from ${\cal T}$. Then, the probability that a randomly selected basic skill is learned can be computed as $\zeta=1-e^{-cR(1-p)}$, where $p$ is the solution to (3). Suppose that the basic skills required by a domain-specific skill are uniformly selected from the set of basic skills $𝒮\color[rgb]{0,0,0}{\cal S}$. Let $\delta_{\ell}$ be the probability that the number of basic skills required by a domain-specific skill is $\ell$ and $\Lambda_{f}(x)=\sum_{\ell=0}^{\infty}\delta_{\ell}x^{\ell}$ be the corresponding generating function. Then we say a domain-specific skill is learnable if all its basic skills are learned in the basic model. Hence, it occurs with probability
$\sum_{\ell=0}^{\infty}\delta_{\ell}{\zeta}^{\ell}=\Lambda_{f}(\zeta).$ (19)
To fine-tune the basic model, we collect a subset of domain-specific texts ${\cal D}_{f}$ and present these $|{\cal D}_{f}|$ texts to a 1-skill learner to acquire domain-specific skills. Let $R_{f}=|{\cal D}_{f}|/|S_{f}|$ be the ratio of the number of training domain-specific texts to the number of domain-specific skills. Similar to the basic model, we make the following two assumptions about sampled domain-specific texts:
(Poisson degree distribution) Denote by $N_{f}(t)$ the number of domain-specific skills in the domain-specific text $t\in{\cal D}_{f}$. Then $\{N_{f}(t),t\in{\cal D}_{f}\}$ are independent Poisson random variables with mean $c_{f}$, i.e., ${\bf\sf P}(N_{f}(t)=k)=e^{-c_{f}}\frac{c_{f}^{k}}{k!}$ for $k=0,1,2,\ldots$.
(Uniform connection) Each edge of a domain-specific text node is connected to a domain-specific skill node independently and uniformly. Thus, the probability that an edge of the domain-specific text node $t$ is connected to a particular domain-specific skill node $s$ is $1/|{\cal S}_{f}|$.
A domain-specific text node is said to be learnable if all of the domain-specific skill nodes connected to it are learnable. Otherwise, it is called a non-learnable domain-specific text node. We assume that nothing can be learned from a non-learnable domain-specific text when it is presented to a learner. From (A1d), (A2d) and (19), we deduce that the probability a randomly selected domain-specific text is learnable is given by
$\sum_{\ell=0}^{\infty}e^{-c_{f}}\frac{c_{f}^{\ell}}{\ell!}(\Lambda_{f}(\zeta))^{\ell}=e^{-c_{f}(1-\Lambda_{f}(\zeta))}.$ (20)
Define the bipartite graph formed by the learnable domain-specific skill nodes and the learnable domain-specific text nodes as the learnable domain-specific bipartite graph. In such a bipartite graph, there are approximately $\Lambda_{f}(\zeta)|{\cal S}_{f}|$ learnable domain-specific skill nodes and $e^{-c_{f}(1-\Lambda_{f}(\zeta))}|{\cal D}_{f}|$ learnable domain-specific text nodes. Furthermore, from Bayes’ formula, we deduce that the degree of a randomly selected learnable domain-specific text node $t$ in the learnable domain-specific bipartite graph follows a Poisson random variable with mean $c_{f}\Lambda_{f}(\zeta)$ since
@@
$=\frac{e^{-c_{f}}\frac{c_{f}^{\ell}}{\ell!}(\Lambda_{f}(\zeta))^{\ell}}{e^{-c_{f}(1-\Lambda_{f}(\zeta))}}=e^{-c_{f}\Lambda_{f}(\zeta)}\frac{(c_{f}\Lambda_{f}(\zeta))^{\ell}}{\ell!}$ (21)
This leads to the following proposition:
@@
Consider the learnable domain-specific bipartite graph. Let $\tilde{\cal S}_{f}$ (resp. $\tilde{\cal D}_{f}$) be the set of learnable domain-specific skills (resp. texts). The learnable domain-specific bipartite graph satisfies the following properties:
(Poisson degree distribution) Denote by $\tilde{N}_{f}(t)$ the number of learnable domain-specific skills in the learnable domain-specific text $t\in\tilde{\cal D}_{f}$. Then $\{\tilde{N}_{f}(t),t\in\tilde{\cal D}_{f}\}$ are independent Poisson random variables with mean $c_{f}\Lambda_{f}(\zeta)$.
(Uniform connection) Each edge of a learnable domain-specific text node is connected to a learnable domain-specific skill node independently and uniformly. Thus, the probability that an edge of the learnable domain-specific text node $t$ is connected to a particular learnable domain-specific skill node $s$ is $1/|\tilde{\cal S}_{f}|$.
In light of Proposition 6, we can apply the same density evolution analysis used in Section III to analyze the learnable domain-specific bipartite graph. Suppose that we still use the 1-skill learner for fine tuning. Let $p_{f}$ be the probability that the text end of a randomly selected edge in the learnable domain-specific bipartite graph is not learned after fine tuning. Note that $|\tilde{\cal S}_{f}|\approx\Lambda_{f}(\zeta)|{\cal S}_{f}|$ and $|\tilde{\cal D}_{f}|\approx e^{-c_{f}(1-\Lambda_{f}(\zeta))}|{\cal D}_{f}|$. Replacing $c$ with $c_{f}\Lambda_{f}(\zeta)$ and $R$ with $R_{f}e^{-c_{f}(1-\Lambda_{f}(\zeta))}/\Lambda_{f}(\zeta)$ in (3) yields
$p_{f}=1-\exp\Big{(}{-c_{f}\Lambda_{f}(\zeta)e^{-(1-p_{f})c_{f}R_{f}e^{-c_{f}(1-\Lambda_{f}(\zeta))}}}\Big{)}.$ (22)
Similarly, let $\zeta_{f}$ be the probability that a randomly selected domain-specific skill node in the learnable domain-specific bipartite graph is learned after fine tuning. Then (cf. (4))
$\zeta_{f}=1-e^{-(1-p_{f})c_{f}R_{f}e^{-c_{f}(1-\Lambda_{f}(\zeta))}}.$ (23)
Now we derive the probability of the testing error of a randomly selected text from the domain-specific texts. Denote such a probability as $\epsilon_{f}$. Assume that this randomly selected domain-specific text has the same distribution as the training data, i.e., it satisfies (A1d) and (A2d). The testing error consists of two cases: (i) a randomly selected text from the set of domain-specific texts ${\cal T}_{f}$ is not learnable, and (ii) a randomly selected text from the set of domain-specific texts ${\cal T}_{f}$ is learnable but one of the domain-specific skills in that text is not learned. As shown in (20), the probability that a randomly selected text from the domain-specific texts is not learnable is $1-e^{-c_{f}(1-\Lambda_{f}(\zeta))}$. Since the degree distribution of a domain-specific text in the learnable domain-specific bipartite graph is Poisson with mean $c_{f}\Lambda_{f}(\zeta)$ in (21), the probability that one of the domain-specific skills in a learnable domain-specific text is not learned is (cf. (5))
$1-\sum^{\infty}_{k=0}e^{-c_{f}\Lambda_{f}(\zeta)}\frac{(c_{f}\Lambda_{f}(\zeta))^{k}}{k!}(\zeta_{f})^{k}$
$=1-e^{-c_{f}\Lambda_{f}(\zeta)(1-\zeta_{f})}.$ (24)
@@
$(1-e^{-c_{f}\Lambda_{f}(\zeta)(1-\zeta_{f})})e^{-c_{f}(1-\Lambda_{f}(\zeta))}.$ (25)
Combining Cases (i) and (ii) in (V) and (25) leads to
$\epsilon_{f}$ $=$ $1-e^{-c_{f}(1-\Lambda_{f}(\zeta))}$ (26)
$+(1-e^{-c_{f}\Lambda_{f}(\zeta)(1-\zeta_{f})})e^{-c_{f}(1-\Lambda_{f}(\zeta))}.$
@@
Figure 11: The contour plot of the probability of the testing error $\epsilon_{f}$ for the 1-skill learner as a function of $R$ and $R_{f}$ with values $c=3$ and $c_{f}=3$.